OpenStudy (anonymous):

integrate by parts: ∫arctan4tdt

OpenStudy (turingtest):

\[u=\tan^{-1}(4t)\]\[dv=tdt\]

OpenStudy (anonymous):

i think you saw an extra t in there dv should just be dt

OpenStudy (turingtest):

yes I did :)

OpenStudy (turingtest):

\[u=\tan^{-1}(4t)\]\[dv=dt\]I'm tired now, g'night!

OpenStudy (anonymous):

\[du = \frac{ 4 }{ 1+t^2 }dt\] \[v = t\]

OpenStudy (anonymous):

\[I = t*\arctan(4t) - 4\int\limits_{}^{}\frac{ t }{ 1+t^2 }dt\] u-sub from here should do it

OpenStudy (fellowroot):